# DiCeption: A Generalist Diffusion Model for Visual Perceptual Tasks

### v2

@@ 5 Conclusion
We have introduced DiCeption, a multi-task visual generalist model based on the diffusion model. Our approach unifies different tasks in the RGB space, leveraging the prior knowledge of pre-trained image generation model to achieve results that are on par with or acceptably lower than the results of specialized foundation models. We achieve good performance without carefully cherry-picking extremely high-quality data or by using an exceptionally large amount of data.
Furthermore, for segmentation tasks, we demonstrate that the strategy of assigning random colors to different instances is highly effective in our framework, enabling high-quality entity segmentation and semantic segmentation. In few-shot fine-tuning, we are able to achieve high-quality results with minimal data and minimal trainable parameters. We believe that DiCeption sheds light on how to effectively use generative model priors to build a strong visual generalist, enabling more efficient solutions to perception tasks.
@@ Appendix A Dataset
Figure S1: Additional visualizations. Our one single model tackles multiple perception tasks.
We summarize the datasets used in our work in Table S1. The depth and normal data samples are obtained by randomly selecting 500K images from OpenImages [51] and labeling them using Depth Pro [7] and StableNormal [123], respectively. The 400K point segmentation data samples are obtained by randomly selecting images from the SA-1B dataset [49]. For the synthesis of point segmentation data, we extract the foreground from P3M-10K [55], AIM500 [57] and AM2K [56], randomly applying transformations such as rotation, resizing, and flipping. These transformed foregrounds are then pasted onto different background images, resulting in 200K synthetic images with fine-grained hair segmentation.
For the validation set, we evaluate depth using the same evaluation protocol as Genpercept [114], conducting tests on the NYUv2 [74], KITTI [32], ScanNet [23], DIODE [102], ETH3D [92]. Similarly, for normal estimation, we followed the evaluation protocol of StableNormal [123] and performed evaluations on the NYUv2 [74], ScanNet [23], DIODE [102]. For point segmentation, we conducted extensive comparisons across 23 datasets. The remaining tasks, including Entity Segmentation, Semantic Segmentation, and Human Keypoints, were evaluated on the MS COCO 2017 dataset [62]. We believe the comprehensive experiments provide solid evidence of the remarkable performance of our method.
@@
Depth 500K OpenImages [51] + Depth Pro [7]
Normal 500K OpenImages [51] + StableNormal [123]
Point Segmentation 400K SA-1B [49]
Point Segmentation 200K P3M-10K [55], AIM500 [57] and AM2K [56]
Human Pose 42K MS COCO 2017 [62]
Semantic Segmentation 120K COCO-Rem [94]
Entity Segmentation 32K EntityV2 [81]
@@
Depth NYUv2 [74], KITTI [32], ScanNet [23], DIODE [102], ETH3D [92]
Normal NYUv2 [74], ScanNet [23], DIODE [102]
Point Segmentation PPDLS [71], DOORS [79], TimberSeg [30], NDD20 [101]
STREETS [95], iShape [120], ADE20K [134], OVIS [80]
Plittersdorf [37], EgoHOS [131], IBD [15], WoodScape [127]
TrashCan [41], GTEA [29, 60], NDISPark [20, 19], VISOR [24, 25]
LVIS [36], Hypersim [90], Cityscapes [22], DRAM [21]
BBBC038v1 [10], ZeroWaste [5], PIDRay [103]
Entity Segmentation MS COCO 2017 [62]
Semantic Segmentation MS COCO 2017 [62]
Human Keypoints MS COCO 2017 [62]
Appendix B Post-processing
@@
Figure S2: Segmentation results on furry objects.
@@
1:  $\mathbf{x}^{\prime}=\operatorname{ExtractRedRegions}(\mathbf{x},\,(255,0,0),\,\sigma)$
2:  $\mathbf{x}_{c}=\operatorname{GetConnectedComponents}(\mathbf{x}^{\prime})$
3:  $\mathbb{C}=\operatorname{GetCircular}(\mathbf{x}_{c})$
4:   $\mathbb{K}_{pred}=\varnothing$
@@
22:  $\mathbb{M}_{pred}=\operatorname{RemoveDuplicateMasks}(\mathbb{M}_{pred},\beta)$
@@
B.1 Post-processing for Keypoints

### v3

@@ 5 Conclusion
We have introduced DiCeption, a multi-task visual generalist model based on the diffusion model. Our approach unifies different tasks in the RGB space, leveraging the prior knowledge of pre-trained image generation model to achieve results that are on par with specialized foundation models. We achieve good performance without carefully cherry-picking extremely high-quality data or by using an exceptionally large amount of data. In few-shot fine-tuning, we are able to achieve high-quality results with minimal data and minimal trainable parameters.
Furthermore, we provide in-depth experimental analyses of strategies for transferring diffusion models to perception tasks. We also discuss the contributions of classifier-free guidance in enhancing model performance, demonstrate that there is no performance gap between our single-task and multi-task model, and highlight the improved detail preservation achieved through pixel-aligned perception training. We believe that DiCeption sheds light on how to effectively use priors of diffusion model to build a strong visual generalist.
@@ Appendix A Dataset
We summarize the datasets used in our work in Table S1. The depth and normal data samples are obtained by randomly selecting 500K images from OpenImages kuznetsova2020open and labeling them using Depth Pro bochkovskii2024depth and StableNormal ye2024stablenormal , respectively. The 400K point segmentation data samples are obtained by randomly selecting images from the SA-1B dataset kirillov2023segment . For the synthesis of point segmentation data, we extract the foreground from P3M-10K li2021privacy , AIM500 li2021deep and AM2K li2022bridging , randomly applying transformations such as rotation, resizing, and flipping. These transformed foregrounds are then pasted onto different background images, resulting in 200K synthetic images with fine-grained hair segmentation.
@@
Depth 500K OpenImages kuznetsova2020open + Depth Pro bochkovskii2024depth
Normal 500K OpenImages kuznetsova2020open + StableNormal ye2024stablenormal
Point Segmentation 400K SA-1B kirillov2023segment
Point Segmentation 200K P3M-10K li2021privacy , AIM500 li2021deep and AM2K li2022bridging
Human Pose 42K MS COCO 2017 lin2015microsoftcococommonobjects
Semantic Segmentation 120K COCO-Rem singh2024benchmarkingobjectdetectorscoco
Entity Segmentation 32K EntityV2 qi2022high
@@
Depth NYUv2 nyu , KITTI kitti , ScanNet scannet , DIODE diode , ETH3D eth3d
Normal NYUv2 nyu , ScanNet scannet , DIODE diode
Point Segmentation PPDLS ppdls , DOORS doors , TimberSeg timberseg , NDD20 ndd20
STREETS streets , iShape ishape , ADE20K ade20k , OVIS ovis
Plittersdorf Plittersdorf , EgoHOS egohos , IBD ibd , WoodScape woodscape
TrashCan trashcan , GTEA gtea ; gtea2 , NDISPark ndispark ; ndispark2 , VISOR visor ; visor2
LVIS lvis , Hypersim hypersim , Cityscapes cityscapes , DRAM dram
BBBC038v1 BBBC038v1 , ZeroWaste zerowaste , PIDRay pidray
Entity Segmentation MS COCO 2017 lin2015microsoftcococommonobjects
Semantic Segmentation MS COCO 2017 lin2015microsoftcococommonobjects
Human Keypoints MS COCO 2017 lin2015microsoftcococommonobjects
For the validation set, we evaluate depth using the same evaluation protocol as Genpercept xu2024diffusion , conducting tests on the NYUv2 nyu , KITTI kitti , ScanNet scannet , DIODE diode , ETH3D eth3d . Similarly, for normal estimation, we follow the evaluation protocol of StableNormal ye2024stablenormal and perform evaluations on the NYUv2 nyu , ScanNet scannet , DIODE diode . For interactive segmentation, we conduct extensive comparisons across 23 datasets. The remaining tasks, including Entity Segmentation, Instance Segmentation, and Human Keypoints, are evaluated on the MS COCO 2017 dataset lin2015microsoftcococommonobjects . We believe the comprehensive experiments on over 30 datasets in total provide solid evidence of the remarkable performance of our method.
Appendix B Additional Analysis
B.1 Token-wise Concat and Channel-wise Concat
We investigated two distinct methodologies for integrating an auxiliary input image into a Diffusion Transformer (DiT) architecture. The first approach involved concatenating the input image tokens with the noisy image tokens along the token dimension, subsequently feeding this combined sequence directly into the DiT model. The second strategy employed channel-wise concatenation of these inputs, followed by a shallow, two-layer Multi-Layer Perceptron (MLP) to align the channel dimensions with the DiT’s input.
Figure S1: Loss curve of token-wise concatenation and channel-wise concatenation.
Constrained by available computational resources, our analysis is conducted within 2 tasks: depth and surface normal estimation. The datasets utilized for depth and surface normal prediction in this ablation are identical to those specified in Table S1. All training hyperparameters remain consistent across both approaches, with the sole architectural divergence being the aforementioned two-layer MLP utilized for feature alignment in the channel-wise concatenation method.
Our findings indicate that the token-wise concatenation strategy is markedly more computationally efficient than its channel-wise counterpart. Specifically, the token-wise approach demonstrates substantially faster convergence speed, as illustrated by the training loss trajectories presented in Figure S1. Furthermore, as demonstrated in Figure S2, channel-wise concatenation is more prone to yielding suboptimal results. We believe that this enhanced efficiency and effectiveness stem from the token-wise concatenation method’s circumvention of additional network parameters. By avoiding the introduction of new trainable components, this strategy appears to more effectively leverage the inherent priors learned by the pre-trained diffusion model. Furthermore, for token-wise concatenation, we independently applied Rotary Position Embeddings (RoPE) su2023roformerenhancedtransformerrotary to both the input image tokens and the noisy tokens. This strategy ensures that corresponding tokens from these two sources share identical positional embeddings, facilitating the model’s rapid learning of their interrelations.
Figure S2: Depth and normal estimation multi-task visualizations comparing channel-wise concatenation, token-wise concatenation, and U-Net are shown. While channel-wise concatenation often leads to suboptimal performance and U-Net struggles with multi-task learning, DiCeption effectively generates high-quality outputs for multiple tasks.
B.2 Architecture of Diffusion Model.
Before the advent of DiT peebles2023scalable , the UNet architecture was predominantly used in diffusion models. We also conduct multi-task experiments based on a UNet pre-trained model SDXL podell2023sdxl . Specifically, we follow Marigold ke2024repurposing by expanding the first convolution layer’s input channels from 4 to 8 to accommodate image inputs, and similarly use task prompts to guide the model in solving different tasks. However, as shown in Figure  S2 and S3 , we find that this approach failed, even for a minimal multi-task scenario involving only depth and normal estimation.
Beyond the established UNet architecture, our research also encompasses an exploration of alternative DiT frameworks, notably PixArt-alpha chen2023pixart , to ascertain the generalizability and efficacy of our proposed methodology when applied to different DiT models. We train DiCeption-PixArt based on the PixArt-alpha-600M model using the same data for training DiCeption and conduct a quantitative evaluation on depth and surface normal prediction, as illustrated in Table S2S3 and S4.
It is pertinent to note that, with a parameter count of approximately 600M, the DiCeption-PixArt variant, while not achieving the same performance benchmarks as our counterpart model trained on the more extensive sd3 architecture, still exhibits a strong capacity for multi-task problem-solving. This multi-tasking proficiency is substantially superior to that of traditional UNet-based models. This result substantiates the versatility of our method and its compatibility with modern transformer-based diffusion models, even with smaller models.
Table S2: Quantitative comparison of depth estimation between ours and ours-PixArt.
Method Training KITTI kitti NYUv2 nyu ScanNet scannet DIODE diode ETH3D eth3d
[…]
Ours 500K 0.069 0.949 0.061 0.960 0.072 0.944 0.289 0.722 0.050 0.975
Ours-PixArt 500K 0.093 0.905 0.096 0.905 0.101 0.901 0.282 0.709 0.071 0.944
Table S3: Quantitative comparison of surface normal estimation between ours and ours-PixArt.
Method Training NYUv2 nyu ScanNet scannet DIODE-indoor diode
[…]
Ours 500K 18.338 10.106 52.850 77.079 82.903 18.842 10.266 53.610 74.895 82.864 16.297 11.117 50.548 83.325 88.774
Ours-PixArt 500K 20.487 12.393 48.663 72.342 80.244 21.663 14.419 37.043 70.781 79.786 17.986 11.190 50.276 79.316 85.248
Table S4: Comparisons of 1-point interactive segmentation between ours and ours-PixArt.
Method Ours-PixArt Ours SAM-vit-h
[…]
Regarding the challenges encountered with UNet-based architectures in multi-task learning paradigms, we posit that their limitations are fundamentally due to two key factors. Firstly, the approach of expanding the input convolution layer introduces additional parameters, thereby potentially disrupting the original model’s inherent prior knowledge. Secondly, the downsampling operations within the U-Net architecture result in a significant loss of information.
Figure S3: The UNet-based model fails to perform multi-task.
B.3 ControlNet
ControlNet zhang2023adding has emerged as a popular approach for integrating novel image conditioning into diffusion models. However, our experiment shows that while ControlNet can learn the general output patterns associated with target tasks, its precision remains notably low, exhibiting limited performance even on single perception task. We train a ControlNet on top of a pre-trained SD3 model for human keypoint estimation. Following the setup of traditional setting zhang2023adding , we introduce ControlNet into the first half of the SD3’s transformer blocks. As depicted in Figure S4, although the model successfully captures the overall visual style of human keypoint predictions, the accuracy of its estimations is significantly deficient.
Figure S4: While ControlNet demonstrates the ability to learn the output modalities of perception tasks, its accuracy remains significantly low. Conversely, our proposed approach yields substantially improved accuracy.
B.4 Classifier-free Guidance
Classifier-free guidance (CFG) ho2022classifierfreediffusionguidance is a technique used in conditional diffusion models to improve the quality of generated samples without additional training. It has become a cornerstone in existing text-to-image models. During inference, it extrapolates from the model’s conditional and unconditional outputs to enhance the influence of the conditioning signal. Specifically, during denoising, the noise at each timestep is a fusion of conditional and unconditional noise:
$n_{t}=n_{t,uncond}+\operatorname{CFG}\cdot(n_{t,cond}-n_{t,uncond}).$ (S1)
Typically, conditional noise $n_{t,cond}$ is the output predicted by the model when conditioned on the prompt embedding, while unconditional noise $n_{t,uncond}$ is the output predicted by the model when conditioned on the negative prompt embedding.
We evaluate the impact of varying CFG values on our multi-task performance. Specifically, our conditional noise $n_{t,cond}$ is the prediction of the model conditioned on the task prompt corresponding to each specific task, while the unconditional noise $n_{t,uncond}$ is the model’s prediction when conditioned on an empty string as the prompt. Our ablation study reveals that a modest application of CFG enhances the quality of depth and normal estimation, yielding perceptibly sharper results. However, this strategy basically has no influence on other tasks such as human keypoints estimation and segmentation, as shown in Figure S5..
Table S5: Interactive Segmentation mIoU of DiCeption across different CFG. CFG has little influence on segmentation
CFG=1 CFG=2 CFG=3 CFG=4 CFG=5
mIoU of 23 Validation Datasets 47.10 47.12 47.08 46.91 46.57
We hypothesize that this is because tasks such as depth and normal estimation inherently demand high precision in the output pixel values to accurately represent continuous geometric surfaces, while other tasks such as human keypoints estimation and segmentation are less sensitive to subtle variations in pixel-level intensities. Additionally, it is also observed that a high CFG scale significantly degrades performance on depth and normal prediction, especially normal prediction. This degradation typically manifests as oversaturated results or the emergence of coarse, granular artifacts, as shown in Figure S5. To further validate our hypothesis, we evaluate the performance of our model across varying CFG values, as presented in the Table S6S7 and S5. The results confirm that a mild CFG scale enhances prediction quality of depth and normal, whereas larger values adversely affect performance.
Table S6: Quantitative comparison of depth estimation with different CFG value.
Method Training KITTI kitti NYUv2 nyu ScanNet scannet DIODE diode ETH3D eth3d
[…]
Ours-CFG=1 500K 0.075 0.945 0.072 0.939 0.075 0.938 0.243 0.741 0.053 0.967
Ours-CFG=2 500K 0.069 0.949 0.061 0.960 0.072 0.944 0.289 0.722 0.050 0.975
Ours-CFG=3 500K 0.092 0.910 0.076 0.938 0.093 0.910 0.343 0.679 0.059 0.966
Ours-CFG=4 500K 0.105 0.876 0.087 0.915 0.104 0.884 0.362 0.654 0.066 0.956
Ours-CFG=5 500K 0.124 0.831 0.097 0.893 0.115 0.863 0.383 0.609 0.072 0.947
Table S7: Quantitative comparison of surface normal estimation with different CFG value.
Method Training NYUv2 nyu ScanNet scannet DIODE-indoor diode
[…]
Ours-CFG=1 500K 18.302 10.538 52.533 75.977 82.573 19.348 12.129 46.410 74.805 82.176 17.946 8.686 62.641 81.152 85.398
Ours-CFG=2 500K 18.338 10.106 52.850 77.079 82.903 18.842 10.266 53.610 74.895 82.864 16.297 11.117 50.548 83.325 88.774
Ours-CFG=3 500K 19.817 10.989 51.312 72.509 79.497 22.287 11.849 49.110 70.075 77.376 18.546 12.475 46.627 76.532 85.398
Ours-CFG=4 500K 21.433 12.012 47.543 69.175 77.003 24.117 13.029 41.334 65.865 73.278 22.886 14.784 41.271 65.661 74.098
Ours-CFG=5 500K 23.352 13.259 43.016 65.727 73.443 26.972 14.364 35.419 57.822 68.776 27.046 19.286 33.349 56.885 66.728
Figure S5: Results on different guidance scale. Depth and normal predictions are highly sensitive to the CFG value, whereas other tasks are barely affected. Based on both the visualization results and the evaluation metrics in Table S6S7 and S5, we set the CFG value to 2 by default.
B.5 Flow-matching Inherently Support Few-step Inference In Perception
We conduct experiments and observe that our model inherently supports few-step inference for perception tasks without any additional techniques, including classifier free guidance, and shows very little performance degradation. The effectiveness of few-step acceleration varies across different tasks. For tasks such as depth and surface normal estimation, the number of inference steps can be reduced to as few as one with acceptable slight performance degradation. For more complex tasks such as interactive segmentation, the model is still able to achieve comparable results using significantly fewer steps while maintaining competitive performance, as demonstrated in Table S8S9 and S10. To the best of our knowledge, this is the first time such a capability is demonstrated in diffusion model for multi-task perception. It strongly supports the advantage of flow-matching-based diffusion models in solving perception tasks.
Table S8: Quantitative comparison of our few-step depth estimation results.
Method Training KITTI kitti NYUv2 nyu ScanNet scannet DIODE diode ETH3D eth3d
[…]
28-step 500K 0.069 0.949 0.061 0.960 0.072 0.944 0.289 0.722 0.050 0.975
14-step 500K 0.077 0.942 0.063 0.958 0.074 0.943 0.272 0.718 0.048 0.978
7-step 500K 0.081 0.939 0.065 0.953 0.078 0.943 0.286 0.714 0.052 0.971
3-step 500K 0.083 0.938 0.069 0.953 0.077 0.940 0.294 0.707 0.063 0.967
1-step 500K 0.086 0.936 0.072 0.945 0.076 0.937 0.305 0.702 0.065 0.967
Table S9: Quantitative comparison of our few-step normal map results.
Method Training NYUv2 nyu ScanNet scannet DIODE-indoor diode
[…]
28-step 500K 18.338 10.106 52.850 77.079 82.903 18.842 10.266 53.610 74.895 82.864 16.297 11.117 50.548 83.325 88.774
14-step 500K 18.631 10.463 52.837 75.288 81.682 18.337 10.579 53.223 75.533 82.631 16.131 11.463 50.849 83.391 88.829
7-step 500K 18.335 10.492 52.771 75.443 81.936 19.008 10.363 52.628 74.886 82.055 16.835 11.330 50.039 82.443 88.218
3-step 500K 18.067 10.417 53.046 76.500 81.673 19.337 10.329 52.223 75.731 82.081 17.205 12.047 50.046 83.010 87.531
1-step 500K 18.094 10.382 51.839 76.575 81.371 19.386 10.395 52.139 75.492 81.879 17.004 11.849 49.808 82.972 87.582
Table S10: Interactive Segmentation mIoU of DiCeption across different inference steps.
28-step 14-step 7-step 3-step 1-step
mIoU of 23 Validation Datasets 47.10 47.01 46.89 45.18 42.53
We believe this is because flow matching explicitly imposes linear constraints at each intermediate denoising step—specifically, each noisy latent is constructed as a linear interpolation between the pure noise and the target signal. This process effectively straightens the denoising trajectory, allowing the model to follow an approximately linear path even when using only a few inference steps. In contrast, if the model is trained solely with one-step denoising, the intermediate steps are not constrained and lacks this linear constraint across the trajectory, thus producing poor results as we show in Section 4.3. In contrast, traditional ODE-based diffusion models do not impose such linear trajectory constraints, and therefore cannot support inference with few denoising steps (such as 4 steps) after being trained with multi-step denoising (such as 50 steps). Our additional experiment proves this. We further experiment with PixArt-alpha chen2023pixart , which uses a DiT-style architecture but adopts a standard ODE-based scheduler. Its results significantly deteriorate when the number of inference steps is reduced, as shown in Table S11, further supporting our analysis.
Table S11: Quantitative comparison of few-step depth estimation results using Pixart-alpha.
Method Training KITTI kitti NYUv2 nyu ScanNet scannet DIODE diode ETH3D eth3d
[…]
20-step 500K 0.093 0.905 0.096 0.905 0.101 0.901 0.282 0.709 0.071 0.944
10-step 500K 0.146 0.872 0.153 0.861 0.159 0.844 0.347 0.658 0.119 0.895
In image generation tasks, simply reducing inference steps in a flow-matching-based text-to-image model also leads to noticeable quality degradation. This is due to the high complexity and variability introduced by diverse text prompts. In contrast, our perception tasks eliminate the influence of textual prompts, which we believe explains why prior works like One Diffusion le2024diffusiongenerate require 50 100 inference steps for denoising while ours works well with just a few steps. For comparisons on inference efficiency, we select One Diffusion as baseline and conduct a comparative study on our shared task, depth estimation, under varying numbers of inference steps, as demonstrated in Table S12. Unlike One Diffusion, which suffers from significant performance degradation during few-step inference and fails to produce reasonable results in the 1-step setting, our method is capable of generating high-quality outputs even with just a single inference step. The results demonstrate that our method significantly outperforms One Diffusion in both efficiency and output quality.
Table S12: Quantitative comparison of One Diffusion and DiCeption in few-step depth estimation. We compared three experimental settings based on the number of steps: the default configuration, a quarter of the default steps, and one single step.
Method KITTI kitti NYUv2 nyu ScanNet scannet DIODE diode ETH3D eth3d
[…]
Ours-28-step 0.069 0.949 0.061 0.960 0.072 0.944 0.289 0.722 0.050 0.975
Ours-7-step 0.081 0.939 0.065 0.953 0.078 0.943 0.286 0.714 0.052 0.971
Ours-1-step 0.086 0.936 0.072 0.945 0.076 0.937 0.305 0.702 0.065 0.967
OD-50-step 0.101 0.908 0.087 0.924 0.094 0.906 0.399 0.661 0.072 0.949
OD-12-step 0.142 0.867 0.114 0.871 0.128 0.853 0.411 0.659 0.092 0.910
OD-1-step FAIL FAIL FAIL FAIL FAIL FAIL FAIL FAIL FAIL FAIL
B.6 Few-shot Finetuning Comparisons on SD3 and Ours
We conduct a comparative evaluation of few-shot tuning performance between SD3 and our DiCeption. All training data and settings are kept identical for both approaches to ensure a fair comparison. Our findings reveal that DiCeption not only adapts to new tasks more rapidly but also achieves better performance post-convergence when compared to SD3. Specifically, Figure S6 (a) illustrates that after convergence, our method yields higher-quality results than SD3 on image highlighting. Furthermore, as depicted in Figure S6 (b), DiCeption demonstrates faster convergence speed. These results collectively underscore the substantial potential of our model for efficient and effective adaptation to novel tasks.
Figure S6: Image highlighting few-shot finetuning comparisons on SD3 and Ours. (a) Our DiCeption achieves better performance. Pixel-level aligned training mitigates generated artifacts. (b) Results of Ours and SD3 in the same training iteration. Our DiCeption is able to adapt to new tasks faster than SD3.
B.7 Pixel-Level Alignment Training Enhances Detail Preservation
We find that training on pixel-level aligned perception tasks endows the model with a strong ability to preserve fine-grained details. We argue that this capability holds significant practical value. For instance, while existing state-of-the-art method IC-Light zhang2025scaling for image relighting can generate visually impressive results, it often suffers from noticeable detail loss such as inconsistency of the individuals’ appearance. In contrast, our approach demonstrates superior fidelity in preserving fine-grained details, including nuances that may not be readily perceptible to the human eye. This is demonstrated in our qualitative comparisons in Figure S7.
It is important to emphasize that our goal is not to compare the lighting quality between methods, but rather to highlight our model’s ability to significantly reduce generative artifacts and retain structural details. We attribute this strength to the model’s exposure to pixel-level aligned tasks during training. Additional comparisons with SD3 esser2024scaling in Figure S6 further support this observation. We consider this finding highly promising and believe it holds substantial implications for detail-preserving generative modeling and downstream applications.
Figure S7: Comparisons of detail preservation, rather than lighting quality. Pixel-level aligned training leads to improved preservation of fine-grained details. Better viewed with zoom-in. Input images are generated and from public available BAID dataset lv2022backlitnet .
Appendix C Post-processing
@@
Figure S8: Segmentation results on furry objects. Our interactive segmentation achieves matting-level accuracy.
@@
[…]
@@
[…]
@@
C.1 Post-processing for Keypoints